# Bohr-Sommerfeld type surgeries and Disk Potential

### v1

@@ Abstract.
We construct a new surgery type operation by switching between two exact fillings of Legendrians which we call a BSP surgery. In certain cases, this surgery can preserve monotonicity of Lagrangians. We prove a wall-crossing type formula for the change of the disk-potential under surgery with Bohr-Sommerfeld profiles. As an application, we show that Biran’s circle-bundle lifts admit a Bohr-Sommerfeld type surgery. We use the wall-crossing theorem about disk-potentials to construct exotic monotone Lagrangian tori in $\mathbb{P}^{n}$.
@@ 1. Introduction
A natural way to construct new manifolds from old is via performing a surgery and gluing operation. In the symplectic world, this idea was utilized in the seminal paper of Polterovich [pol:surg] to construct a surgery type operation on Lagrangian submanifolds which resolved singular points arising at transverse intersections. Subsequently, numerous surgery operations were developed to construct new Lagrangians from old, see e.g. [haug, makwu, PT20, Cha23].
A connecting thread among these surgery operations is that all of them switch between two fillings of certain Legendrians. Given a symplectic manifold $(M,\omega)$ and a separating contact-type hypersurface $Z$, we can split $M$ along $Z$ by viewing $M\setminus Z$ as $M\setminus Z=M^{+}\sqcup M^{-},$ where $M^{+}$ is a capping and $M^{-}$ is a filling of the contact manifold $Z$. If $L$ is a Lagrangian in $M$ such that it intersect the contact-type hypersurface $Z$ at a Legendrian $\Lambda$, then $Z$ splits $L$ similarly, $L\setminus Z=L^{+}\sqcup L^{-}$ where $L^{+}$ is a Lagrangian capping and $L^{-}$ is a Lagrangian filling of $\Lambda$. We can construct a new Lagrangian $\widetilde{L}$ by replacing the filling $L^{-}$ with a different filling $\widetilde{L}^{-}$ and gluing along the Legendrian $\Lambda,$
@@
For example, we can view the Polterovich-surgery as switching between fillings of a Legendrian consisting of two disjoint Legendrian spheres.
Although this idea is simple to state, it is quite challenging to understand how such surgeries affect Floer theoretic invariants. The effect of such surgeries have been carried out in certain special cases such as for Polterovich surgery see [ch10, PW19, seidelFuk], higher mutation see [PT20, Cha23].
In this article, we construct a generalization of Polterovich-surgery and local higher mutations and study their effect on disk-potentials (also known as super-potential). The new surgery operation switches between exact fillings of a Legendrian $\Lambda_{2}$  [^1] The subscript $2$ denotes that there are two connected components in this Legendrian as we will see in §2.2 that is obtained by a Bohr-Sommerfeld lift as in Dimitroglou Rizell-Golovko [rg19]. We prove that such surgery operations preserve monotonicity of Lagrangians under certain assumptions. We call such surgery operations Bohr-Sommerfeld surgery with a $\Lambda$-profile and notate the operation as $BSP^{\Lambda}(L)$. In the article we will drop ${\Lambda}$ from the notation when the Legendrian is clear from context.
The main Floer-theoretic result in this paper is the following change-of-potential formula. This can be viewed as a generalization of the mutation formula for Lagrangian disk-surgery and higher mutations, see [Aur07, PT20, PW19]. We will drop some hypotheses for the sake of clarity of exposition, the curious reader should refer to Theorem 3.7 for the details of the theorem.
{mdframed}
Main Theorem (Theorem 3.7).
Under nice conditions, a Bohr-Sommerfeld profile surgery of $L$ with a $\Lambda$-profile results in the following change of disk potential
where $W_{\Lambda}$ is the augmentation polynomial of $\Lambda$.
@@
By specializing to the case of BSP surgery with a profile modelled on the lift of the equator circle in $\mathbb{P}^{1}$ we can recover the change of disk-potential under Lagrangian disk surgery (Example 3.8) and by specializing to the lift of the Clifford torus we obtain the change of disk potential under higher mutation (Example 3.9).
Remark 1.2.
We expect that the change of potential formula in Theorem 3.7 can be upgraded to an invariance result in the Donaldson-Fukaya category. In particular, under a birational change of local systems, $\rho\mapsto\rho_{\Lambda},$ we expect $(L,\rho)$ and $(BSP(L),\rho)$ to be quasi-isomorphic to each other. The proof strategy presented here using SFT breaking should work, but we do not pursue this direction in this article.
@@ 1.1. Applications
We explore a few applications of BSP surgery before proceeding to the construction. Building on Biran’s circle-bundle construction [Biran_lagrangina_barrier, b06, bc09], we show how Biran lifts can be Hamiltonian isotoped to include BSP handles, allowing for monotone BSP surgeries and thus produce exotic circle bundle lifts. We also show that BSP surgery provides a geometric realization of mutation of potentials as studied in [21coates, cruzgalkin, minksum]. Furthermore, we apply these ideas to construct more exotic monotone Lagrangian tori.
1.1.1. Conical Surgeries
For Lagrangian surfaces which have a Lagrangian disk cleanly intersecting in a circle, one can perform a Lagrangian disk surgery as introduced by [yau17] to get a new Lagrangians. This procedure can be thought of performing an anti-surgery using the Lagrangian disk to create a doubly-immersed-singular point and then resolving it by doing a Polterovich surgery in the ‘opposite’ direction. Such surgeries have been widely used, especially in studying exact Lagrangian fillings, see [stw1, stw2, cginfty, cwmicro].
We define a notion of conical surgery, a generalization of disk-surgery. Given a Bohr-Sommerfeld lift $\Lambda$ in the standard contact sphere $S^{2n+1}$, there is a natural Lagrangian cone $C(\Lambda)$ supported on $\Lambda$. We define a Lagrangian cone modelled on a Bohr-Sommerfeld Legendrian $\Lambda$ to be a topological embedding of the cone $C(\Lambda)$ such that :
  • Away from the conical point, the embedding is actually a smooth Lagrangian embedding.
  • The embedding looks like $C(\Lambda)$ in a Darboux ball, centered at the conical point.
@@ 1.1.2. Exotic Lagrangian Circle-Bundles
Biran’s circle-bundle construction ([b06, §4]) provides a way to lift Lagrangians from symplectic divisors in symplectic manifolds. Later, Biran-Cornea [bc09] proved that there are monotone circle-bundle lifts $\widetilde{L}\hookrightarrow M$ from monotone Lagrangians $L$ in symplectic divisors $D\hookrightarrow M$. The relation between the disk-potential of $\widetilde{L}$ with that of $L$ has been studied in the upcoming work [DTVW] using [bk13].
@@
Thus, we can apply Theorem Main Theorem to see that the disk potential
(1.1.1) $W_{BSP^{\Lambda}(\widetilde{L})}(\cdots)=W_{\widetilde{L}}(\cdots,zW_{\Lambda},\cdots)$
where $z$ is defined as the monomial corresponding to a cycle which passes through the BSP handle once.
@@
$W_{\Lambda}$ is defined as a lift of the disk potential $W_{L}$ under the natural pullback map induced from $\Lambda\rightarrow L$. Just from knowing that disk-potentials are Laurent polynomials, Equation 1.1.1 indicates that the disk potential of the circle bundle lift “sees” that of the base. In particular, the potential of the circle bundle lift admits a change of variables such that swapping $z$ with $zW_{\Lambda}$ keeps the resulting rational map a Laurent polynomial. This, of course, is evident clearly from the disk potential computation of the Biran circle-bundle Lagrangian in [bk13, DTVW].
Recall that Vianna [Via16, Via17] proved that there are infinitely many monotone Lagrangian tori in $\mathbb{C}P^{2}$. In fact, Vianna proves that for every Markov triple $(a,b,c)$ 222 A Markov triple $(a,b,c)$ is a positive integer solution to the Markov equation $a^{2}+b^{2}+c^{2}=3abc$. Later in [DTVW, chw] it was proved that the Biran lifts of the Vianna tori remain in distinct symplectomorphism classes. We will show that performing BSP surgery on the (iterated) Biran lifts of the Vianna tori will produce exotic monotone tori which have not yet appeared in the literature to the best of our knowledge.
@@ 1.1.3. Geometric realization of mutations of polynomials
Mutations of Laurent polynomial have turned to be quite essential in mirror symmetry, see e.g. [cruzgalkin, 21coates, minksum]. The effect of BSP surgery on the disk-potential can be phrased in terms of such mutations, in particular BSP surgery on a Lagrangian corresponds to performing certain combinatorial mutation of a Laurent polynomial.
@@
A Newton polytope is defined to be a Fano polytope ( see [minksum] ), if
@@
$width_{w}(P)=h_{max}-h_{min}.$
We define slices $P_{i}$ as
$P_{i}=\text{conv}\{a\in P|w(a)=i\}.$
Note that $P_{i}=\emptyset$ for $i<h_{min}$ and $i>h_{max}$. Assume that there is a convex lattice polytope $F\subset w^{\perp}$ 333 $w^{\perp}$ is the subspace $\{p|w(p)=0\}$., such that for every height $h_{min}\leq h<0$ there exists a (possibly empty) lattice polytope $G_{h}$ satisfying
@@
where $+$ denotes Minkowski sum. We call such an $F$ a factor for $P$ with respect to $w$. We define the combinatorial mutation given by width vector $w$, factor $F$, and polytopes $G_{h}$ to be the convex lattice polytope
@@
Although mutation uses the existence of polytopes $G_{h}$, from Proposition 1 [minksum] one has that the mutation is independent of the choice of $G_{h}$. If we write the disk potential $W_{L}$ of a monotone Lagrangian with BSP-handle as a polynomial
@@
where $k<0,$ and $l>0$, $C_{i}$ are Laurent polynomials with variable $x_{i},y_{j}$. Particularly in the case of Biran circle bundle lifts, the newton polytope changes by a combinatorial mutation with a factor corresponding to a codimension one face. From Theorem Main Theorem we can see that $W_{\Lambda}^{-h}$ divides $C_{h}$ for all $k\leq h<0$.
@@
where $\hat{z}$ is the dual of the $z$ variable. Thus, we see that the change of disk-potential in Theorem Main Theorem amounts to a combinatorial mutation with a factor corresponding to $W_{\Lambda}$.
@@ 1.1.4. Exotic tori in $\mathbb{P}^{n}$
To see that the exotic lifts of the Vianna tori using BSP surgery are, in fact, different from Biran lifts of Vianna tori, we study the effect of BSP surgery on the Newton polytope of the disk potential using combinatorial mutations.
Note that the width of $\operatorname{Newt}(W_{\overline{T_{(a,b,c)}}})$ is 4 and $h_{max}=3$. Thus, from the combinatorial mutation formulation, we see that the Newton polytope of $BSP^{\Lambda_{(a,b,c)}}(\overline{T}_{(a,b,c)})$ has a face isomorphic to $3\operatorname{Newt}(T_{(a,b,c)})$. Here $3\operatorname{Newt}(T_{(a,b,c)})$ refers to the triple Minkowski sum, $\operatorname{Newt}(T_{(a,b,c)})+\operatorname{Newt}(T_{(a,b,c)})+\operatorname{Newt}(T_{(a,b,c)})$.
@@
Under a judicious choice of basis of $H_{1}(\overline{T}^{3}_{(1,1,2)})$ we have
@@
and thus, from our wall-crossing formula we have

### v2

# arXiv:2409.11603v2[math.SG] 27 Feb 2026

@@ Abstract.
We construct a new surgery type operation for Lagrangian submanifolds by switching between two exact fillings of Legendrians which we call a BSP surgery. In certain cases, this surgery can preserve monotonicity of Lagrangians. We prove a wall-crossing type formula for the change of the disk-potential under surgery with Bohr-Sommerfeld profiles. As an application, we show that Biran’s circle-bundle lifts admit a Bohr-Sommerfeld type surgery. We use the wall-crossing theorem about disk-potentials to construct exotic monotone Lagrangian tori in $\mathbb{P}^{n}$. Moreover, we use Bohr-Sommerfeld surgeries to generalize the notion of Lagrangian disk surgeries.
@@ 1. Introduction
Surgeries provide a natural way to construct new manifolds from old ones. In the symplectic world, this idea was introduced in the pioneering paper of Polterovich [pol:surg]. Subsequently, numerous Lagrangian surgery operations were developed using Polterovich surgery as the main framework, see e.g. [yau17, haug, makwu]. Recently, a new surgery operation, local higher mutation, was introduced in [Cha23] that resolved certain conical singularities modeled on the Clifford torus. The construction of local higher mutation was inspired by higher mutations of Lagrangian tori introduced in [PT20].
A connecting thread among these surgery operations is that all of them switch between two fillings of certain Legendrians. Given a symplectic manifold $(M,\omega)$ and a separating contact-type hypersurface $Z$, we can split $M$ along $Z$ by viewing $M\setminus Z$ as $M\setminus Z=M^{+}\sqcup M^{-},$ where $M^{+}$ is a capping and $M^{-}$ is a filling of the contact manifold $Z$. If $L$ is a Lagrangian in $M$ such that it intersects the contact-type hypersurface $Z$ at a Legendrian $\Lambda$, then $Z$ splits $L$ similarly, $L\setminus Z=L^{+}\sqcup L^{-}$ where $L^{+}$ is a Lagrangian capping and $L^{-}$ is a Lagrangian filling of $\Lambda$. We can construct a new Lagrangian $\widetilde{L}$ by replacing the filling $L^{-}$ with a different filling $\widetilde{L}^{-}$ and gluing along the Legendrian $\Lambda,$
@@
Although this idea is simple to state, it is quite challenging to understand how such surgeries affect Floer theoretic invariants. The Floer theoretic aspects of such surgeries have been studied in certain special cases, such as for Polterovich surgery [ch10, PW19, seidelFuk], higher mutation [PT20, Cha23].
In this article, we construct a large class of surgery type operations which generalize Polterovich-surgery and local higher mutations and study the effect of such surgeries on disk-potentials. More precisely, for every Bohr-Sommerfeld Legendrian $\Lambda,$ as introduced in [rg19], we construct a surgery operation $BSP^{\Lambda}$. The new surgery operation switches between exact fillings of a Legendrian $\Lambda_{2}$  [^1] The subscript $2$ denotes that there are two connected components in this Legendrian as we will see in §2.2 that is obtained from ‘doubling’ a Bohr-Sommerfeld Legendrian $\Lambda$. We prove that such surgery operation preserves monotonicity of Lagrangians under certain assumptions. We call such a surgery operation Bohr-Sommerfeld surgery with a $\Lambda$-profile and denote the surgered Lagrangian as $BSP^{\Lambda}(L)$. In the article, we will drop $\Lambda$} from the notation when the Legendrian is clear from context.
The main Floer-theoretic result in this paper is a wall-crossing style formula. This can be viewed as a generalization of the mutation formula for Lagrangian disk-surgery and higher mutations, [Aur07, PT20, PW19]. We will drop some hypotheses for the sake of clarity of exposition; the curious reader should refer to Theorem 3.7 for details of the theorem.
Main Theorem 1 (Theorem 3.7).
Under nice conditions, a Bohr-Sommerfeld profile surgery of $L$ with a $\Lambda$-profile results in the following change of disk potential $W_{BSP^{\Lambda}(L)}(x_{1},\dots,x_{k},z,w_{1},\dots,w_{l})=W_{L}(x_{1},\dots,x_{k},zW_{\Lambda}(x_{1},\dots,x_{k}),w_{1},\dots,w_{l}).$ where $W_{\Lambda}$ is the augmentation polynomial of $\Lambda$.
Remark $1.$
(Lagrangian Mutations as a special case of BSP surgery) By specializing to BSP surgery with a profile modeled on the real Legendrian circle $S^{1}_{\mathbb{R}}=S^{3}\cap\mathbb{R}^{2}$ in the standard contact sphere $S^{3}\hookrightarrow\mathbb{C}^{2}$, we can recover the classical wall-crossing formula [Aur07, PW19, ch10] for Lagrangian disk surgeries (Example 3.8). By specializing to the BSP surgery, corresponding to the Legendrian Clifford torus, we obtain the change of disk potential under a local higher mutation (Example 3.9).
@@
We expect that the change of potential formula in Theorem 3.7 can be upgraded to an invariance result in the Donaldson-Fukaya category. In particular, under a birational change of local systems, $\rho\mapsto\rho_{\Lambda},$ we expect $(L,\rho)$ and $(BSP^{\Lambda}(L),\rho)$ to be quasi-isomorphic to each other. The proof strategy presented here using SFT breaking should work, but we do not pursue this direction in this article.
@@ 1.1. Applications
We explore a few applications of BSP surgery before proceeding to the construction. Building on Biran’s circle-bundle construction [Biran_lagrangina_barrier, b06, bc09], we show how Biran lifts can be Hamiltonian isotoped to include BSP handles, allowing for monotone BSP surgeries and thus producing exotic circle bundle lifts. We also show that BSP surgery provides a geometric realization of mutation of potentials as studied in [21coates, cruzgalkin, minksum]. Furthermore, we apply these ideas to construct exotic monotone Lagrangian tori in $\mathbb{C}P^{n}$.
1.1.1. Lagrangian Cone Surgeries
For Lagrangian surfaces that have a Lagrangian disk cleanly intersecting in a circle, one can perform a Lagrangian disk surgery as introduced by [yau17] to get a new Lagrangians. This procedure can be thought of as performing an anti-surgery using the Lagrangian disk to create a doubly-immersed-singular point and then resolving it by doing a Polterovich surgery in the ‘opposite’ direction. Such surgeries have been widely used, especially in studying exact Lagrangian fillings [stw1, stw2, cginfty, cwmicro].
We define the notion of Lagrangian cone surgery, a generalization of disk-surgery : Given a Bohr-Sommerfeld lift $\Lambda$ in the standard contact sphere $S^{2n+1}$, there is a natural Lagrangian cone $C(\Lambda)$ supported on $\Lambda$. We define a Lagrangian cone modeled on a Bohr-Sommerfeld Legendrian $\Lambda$ to be a topological embedding of the cone $C(\Lambda)\cong\Lambda\times[0,1]/(\Lambda\times\{0\})$ such that :
  1. (C.1)
    away from the conical-singularity point, the embedding is a smooth Lagrangian embedding,
  2. (C.2)
    in a Darboux chart centered at the conical point, the embedding looks like $C(\Lambda)$.
In §6.2 we generalize Lagrangian disk surgeries by using BSP surgeries with profiles coming from Legendrians in the standard contact sphere. In particular, we show the following result that proves the existence of Lagrangian cones satisfying properties C.1 and C.2, ensures that a Bohr-Sommerfeld surgery can be performed,
Proposition.
Let $L$ be a monotone Lagrangian and $C$ be a Lagrangian cone modeled on a Bohr-Sommerfeld Legendrian $\Lambda$ in the standard contact sphere $S^{2n+1}$. If $C$ intersects $L$ only at its boundary $\partial C$, and the intersection is clean, then $L$ admits a Bohr-Sommerfeld-Profile surgery modeled on $\Lambda$ such that the surgered manifold $BSP^{\Lambda}(L)$ is also monotone.
@@ 1.1.2. Exotic Lagrangian Circle-Bundles
Biran’s circle-bundle construction ([b06, §4]) provides a way to lift Lagrangians from symplectic divisors in symplectic manifolds. Later, Biran-Cornea [bc09] proved that there are monotone circle-bundle lifts $\widetilde{L}\hookrightarrow M$ from monotone Lagrangians $L$ in symplectic divisors $D\hookrightarrow M$. The relation between the disk-potential of $\widetilde{L}$ and that of $L$ has been studied in the upcoming work [DTVW] using [bk13].
@@
Thus, we can apply Theorem 1 to observe the following relationship between the disk potentials,
(1.1.1) $W_{BSP^{\Lambda}(\widetilde{L})}(\cdots)=W_{\widetilde{L}}(\cdots,zW_{\Lambda},\cdots),$
where $z$ is defined as the monomial corresponding to a cycle that passes through the BSP handle once.
@@
$W_{\Lambda}$ is defined as a lift of the disk potential $W_{L}$ under the natural pullback map induced from $\Lambda\rightarrow L$. Just from knowing that disk-potentials are Laurent polynomials, Equation 1.1.1 indicates that the disk potential of the circle bundle lift “sees” that of the base. In particular, the potential of the circle bundle lift admits a change of variables such that swapping $z$ with $zW_{\Lambda}$ keeps the resulting rational map a Laurent polynomial. This, of course, is evident from the disk potential computation of the Biran circle-bundle Lagrangian in [bk13, DTVW].
Recall that Vianna [Via16, Via17] proved that there are infinitely many monotone Lagrangian tori in $\mathbb{C}P^{2}$. In fact, Vianna proved that for every Markov triple [^2] A Markov triple $(a,b,c)$ is a positive integer solution to the Markov equation $a^{2}+b^{2}+c^{2}=3abc$. $(a,b,c)$, there is a monotone Lagrangian torus $T_{(a,b,c)}$. Further, Vianna showed there are no symplectomorphisms of $\mathbb{C}P^{2}$ which map the torus $T_{(a,b,c)}$ to the torus $T_{(a^{\prime},b^{\prime},c^{\prime})}$ corresponding to different Markov triple $(a^{\prime},b^{\prime},c^{\prime})$ . Later in [DTVW, chw] it was proved that the Biran lifts of the Vianna tori remain in distinct symplectomorphism classes. We will show that performing BSP surgery on the (iterated) Biran lifts of the Vianna tori will produce exotic monotone tori which have not yet appeared in the literature to the best of our knowledge.
@@ 1.1.3. Geometric realization of mutations of polynomials
Mutations of Laurent polynomial have turned out to be quite essential in mirror symmetry, e.g. [cruzgalkin, 21coates, minksum]. The effect of BSP surgery on the disk-potential can be phrased in terms of such mutations, in particular BSP surgery on a Lagrangian corresponds to performing certain combinatorial mutations of a Laurent polynomial.
@@
A Newton polytope is defined to be a Fano polytope ( [minksum] ), if
@@
In the equations above, $w(v)$ should be interpreted as $\langle w,v\rangle$ where $\langle\;,\;\rangle$ is the usual inner-product on the Euclidean space. The $w$-width of the polytope $P$ is defined to
$width_{w}(P)=h_{max}-h_{min}.$
We define slices $P_{i}$ for integers $i$ as
$P_{i}=\text{conv}\{a\in P|w(a)=i\}.$
@@
where ++ denotes Minkowski sum. We call such an $F$, a factor for $P$ with respect to $w$. We define the combinatorial mutation given by width vector $w$, factor $F$, and polytopes $G_{h}$ to be the convex lattice polytope
@@
Although mutation uses the existence of polytopes $G_{h}$, Proposition 1 [minksum] proves that the mutation is independent of the choice of $G_{h}$. The effect of the mutation formula in Theorem 3.7 can be studied clearly when we write the disk potential $W_{L}$ of a monotone Lagrangian as a polynomial
@@
where $k<0,$ and $l>0$, $C_{i}$ are Laurent polynomials with variable $x_{i},y_{j}$. Particularly in the case of Biran circle bundle lifts, the Newton polytope changes by a combinatorial mutation with a factor corresponding to a codimension one face. From Theorem 1 we can see that $W_{\Lambda}^{-h}$ divides $C_{h}$ for all $k\leq h<0$.
@@
where $\hat{z}$ is the dual of the $z$ variable. Thus, we see that the change of disk-potential in Theorem 1 amounts to a combinatorial mutation with a factor corresponding to $W_{\Lambda}$.
@@ 1.1.4. Exotic tori in $\mathbb{P}^{n}$
To see that the exotic lifts of the Vianna tori using BSP surgery are, in fact, different from the Biran lifts of Vianna tori, we study their Newton polytopes. More precisely, we study the effect of BSP surgery on the Newton polytope of the disk potential using combinatorial mutations.
From [chw] we know that $\operatorname{Newt}(\overline{T_{(a,b,c)}})$ can be viewed as a cone over $\operatorname{Newt}(T_{(a,b,c)})$. Here $T_{(a,b,c)}$ is the two dimensional Vianna torus corresponding to the Markov triple (a,b,c){(a,b,c)}. Let $w$ be the primitive integral vector perpendicular to the triangular face $\operatorname{Newt}(T_{(a,b,c)})$ contained in $\operatorname{Newt}(\overline{T_{(a,b,c)}})$. Note that the $w$-width of $\operatorname{Newt}(W_{\overline{T_{(a,b,c)}}})$ is 4 and $h_{max}=3$. Thus, from the combinatorial mutation formulation, we see that the Newton polytope of $BSP^{\Lambda_{(a,b,c)}}(\overline{T}_{(a,b,c)})$ has a face isomorphic to $3\operatorname{Newt}(T_{(a,b,c)})$. Here $3\operatorname{Newt}(T_{(a,b,c)})$ refers to the triple Minkowski sum, $\operatorname{Newt}(T_{(a,b,c)})+\operatorname{Newt}(T_{(a,b,c)})+\operatorname{Newt}(T_{(a,b,c)})$.
@@
[…]
@@
and thus, from our wall-crossing formula, we have